OpenStudy (anonymous):

Let F of x equals the integral from 1 to 3 times x of the natural logarithm of t squared. Use your calculator to find F″(1).

OpenStudy (irishboy123):

\[F(x) = \int\limits_{1}^{3x} \ln (t^2) dt\] \[F''(x) = \frac{d^2F}{dx^2}\] yes?